6. Find two independent Frobenius series solutions of y + 4x²y = 0.

Please give a detailed solution, showing all the steps.

Thanks,
Answer #1

Spletton: The given cquction is Y-48!+ 43pPy = 0-0 let yum, y=mumi, glla mm-1) M-2 Put these values in cano m (m-1), M.2 t mputros therefore tror 4 in last term of en Refore can & colll be eam ar mitrs [[max-1)-1] Mto-2 +8 uagt2 = 0 => È qoymato) Ka